OpenStudy (anonymous):

find the value of tan(7pi/6)

OpenStudy (debbieg):

Hmmm... that's true, but you don't need to do that. Multiples of \(\pi/6\) are common unit circle angles, so you just need to use \(\pi/6\) as your reference angle. \(7\pi/6\) is in which quadrant? What is the sign of tangent, in that quadrant? Then, since \(\pi/6\) is your reference angle, you just need to use the fact that: \(\tan7\pi/6=\pm \tan\pi/6\), with the sign depending on the sign of tangent in the quadrant that you're in.

OpenStudy (praxer):

\[\pi/6 \] is less than \[\pi/2\] so it is in the first quadrant so the value will be positive.

OpenStudy (debbieg):

Wha??? \(\pi/6\) is definitely in the first quadrant. But that is only the REFERENCE angle. \(7\pi/6\) is the angle in question.

OpenStudy (praxer):

Multiply it with the general equation then and find the answer.